Hi! I've been stuck on this problem for a while now. If someone could show me how to do this problem, it would be a great help. Thank you! :)
One number is 2 more than 5 times another. Their product is 16. Find the numbers. (There should be 4 values/answers)

Answer by ankor@dixie-net.com(22740)

Two number; a & b
:
One number is 2 more than 5 times another.
a = 5b + 2
:
Their product is 16.
a * b = 16
Replace a with (5b+2) from the first statement
(5b+2)*b = 16
5b^2 + 2b = 16
A quadratic equation
5b^2 + 2b - 16 = 0
Factors to
(5b - 8 )(b + 2) = 0
Two solution for b
5b = 8
b = 8/5
b = 1.6
and
b = -2
:
Find the numbers. (There should be 4 values/answers)
Find a when b = 1.6
a = 5(1.6) +2
a = 8 + 2
a = 10
Find a when b = -2
a = 5(-2) + 2
a = -10 + 2
a = -8
:
:
Check both solutions see if their products = 16
10 * 1.6 = 16
and
-8 * -2 = 16